PLEASE ANSWER BOTH PARTS OF THE QUESTION. I know answer to first part is $66, what about second part??

Refer to table 1. Consider a basket of goods that includes 10 apples, 5 bananas, 3 kiwis, and 1 pineapple. At the prices given in the table, what is the nominal cost of the basket of goods in 2004? Table 1 ear apple baaa kiwi pineapple 2003 $4 $2 $5 $6 2004 $3 $2 $6 $8 Select one: O a. $80 o b. $66 O c. $73 O d. $40 Refer to table 1 again. Consider the same basket of goods (10 apples, 5 bananas, 3 kiwis, and 1 pineapple). At the prices given in the table, what is the rate of inflation between 2003 and 2004 (Use 2003 as the base year)? Table 1 apple banana kiwilpineapple 2003 $4 $2 $5 $6 2004 $3 $2 $6 $8 Select one: O a. 7% О с. 3% O d.-2%
Answer #1

Answer 1 : Option B is correct.

CPI 2004 = ( price of apple * Goods+ price of banana* Goods+ price of kiwi * Goods + price of pineapple* Goods)

CPI 2004 =( $3*10+ $2* 5+$6*3+ $8*1)= $30+$10+$18+$8= $66

Answer2 : Option B is correct.

CPI 2004 =$ 66

CPI 2003 = ( price of apple* Goods + price of banana* Goods + price of kiwi * Goods + price of pineapple * Goods)

CPI 2003 = ( $4* 10+ $2* 5+ $5*3+ $6*1)= ($40+$10+$15+$6)= $71

Inflation rate= (CPI 2004 -CPI 2003)/CPI 2003 = ( 66-71)/71= -7%